OpenStudy (anonymous):

integrate 1/((x+5)^2(x-1))

OpenStudy (anonymous):

man these are brutal

OpenStudy (anonymous):

another damned partial fraction but this one sucks because of the square in the denominator. you have to write is as \[\frac{a}{x-1}+\frac{b}{x+5}+\frac{cx+d}{(x+5)^2}\]

OpenStudy (anonymous):

i don't wanna do it. do you know how?

OpenStudy (anonymous):

lol, im trying to do them .

OpenStudy (anonymous):

this one stinks you have to add the numerator and set the result = 1

OpenStudy (anonymous):

\[a(x+5)^2+b(x+5)(x-1)+(cx+d)(x-1)=1\] and solve for a, b, c, and d

OpenStudy (anonymous):

you can start by replacing x by 1 to get \[a(1+5)^2=1\] so \[a=\frac{1}{36}\]

OpenStudy (anonymous):

the rest is a pain. i would cheat

OpenStudy (anonymous):

lolol i think that's what im going to do. thanks for ALL your help though :)

OpenStudy (anonymous):

yw. you have to multiply out and equate like coefficients. a real drag. but here http://www.wolframalpha.com/input/?i=integrate+1%2F%28%28x%2B5%29^2%28x-1%29%29